OpenStudy (abby09sj09):

Find an equation for the ellipse that satisfies the given conditions. eccentricity = 1/3 foci = (0, +-7) I thought it would be x^2/392 + y^2/348 = 1 but I've tried several times and keep getting different answers...any help would be appreciated! Thanks :)

OpenStudy (eliesaab):

\[ \frac 7 a=e=\frac 1 3\implies a=21 \]

OpenStudy (eliesaab):

\[ b^2=a^2-c^2= 21^2 -7^2=392 \]

OpenStudy (eliesaab):

\[ a^2=21^2=441 \]

OpenStudy (eliesaab):

\[ \frac {x^2}{441}+\frac {y^2}{392}=1 \]

OpenStudy (abby09sj09):

Thank you so much! :) @eliesaab

OpenStudy (eliesaab):

YW